The Gold King Mine spill released 9.3 acre feet of Acid Mine Drainage (AMD) and an estimated 400 kg of Copper. 1 acre-foot = 1233 m3 Colorado Department of Public Health and the Environment (CDPHE) allowable concentration for dissolved copper Cu+2 = 50 µg/L(a) Assume that all copper was initially in its ionic form (Cu2+, which is not a realistic assumption, as much would be sorbed to the iron(III) hydroxide precipitate evident in the video). After one day, the total volume of river and AMD was estimated to be 1400 acre feet. Assuming no precipitation of Cu+2 occurred and after one day the mixing was uniform, what is concentration of Cu+2 in µg/L?
Answer #1

Total volume of river and AMD = 1400×1233 m3 = 1726200m3

Concentration of Cu2+= 400×1000÷1726200 = 0.231723 g/m3

1m3 = 1000 litre

Concentration of Cu2+ =0.231723÷1000 = 0.000231723g/L

Which is equal to 231.723g/L.